OpenStudy (anonymous):

Select one of the factors of x2 + 11x + 24 I answered (x + 3) would this be correct?

OpenStudy (amistre64):

when x=-3 does the expression equal 0?

jhonyy9 (jhonyy9):

(x+3)(x+8)

OpenStudy (anonymous):

As in (x + 3)(x + 8)

OpenStudy (anonymous):

Ami, that's how I could check it?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Sweet! Didn't realize that.

OpenStudy (amistre64):

if (x+a) is a factor; then when x=-a your expression equals zero

OpenStudy (anonymous):

Wait, so if a is already a negative, what would I do?

OpenStudy (anonymous):

(x + (-a))?

OpenStudy (amistre64):

x+(-a) ; then x=-(-a) = a

OpenStudy (anonymous):

Alright, thanks a bunch!

OpenStudy (amistre64):

whatever x value that makes your factor go zero :)

OpenStudy (amistre64):

x+3 = 0 when x=-3 x-9 = 0 when x=9

OpenStudy (anonymous):

In my case, I would replace the variable "x" with -3? -3^2 + 11(-3) + 24?

OpenStudy (amistre64):

yes (-3)^2 +11(-3) + 24 ?= 0 9 -33 +24 ?= 0 -24 + 24 = 0

OpenStudy (anonymous):

-3^2 = 9 9 - 33 + 24 -33 + 33 = 0

OpenStudy (anonymous):

Yes! :D

OpenStudy (amistre64):

any x value that makes it go zero will be a root; and hence you can construct a factor from it

OpenStudy (anonymous):

Thank you! Factoring is making a lot more sense now.

OpenStudy (amistre64):

:) good luck